# SGDFuse: SAM-Guided Diffusion for High-Fidelity Infrared and Visible Image Fusion

### v2

@@ 4.5.1 Object Detection
Fig. 11 shows that fused images significantly improve detection accuracy over single-modality inputs. Our method outperforms mainstream
fusion approaches in both “Person” and “Car” categories, demonstrating superior object perception and generalization.
@@
As shown in Fig. 12, our method outperforms others in accurately localizing multiple targets by effectively preserving infrared thermal information and integrating visible details, thereby enhancing detection accuracy, robustness, and adaptability in complex scenes.
@@ 4.5.2 Semantic Segmentation
Table V shows segmentation results for six categories (Car, Person, Bike, Color Cone, Background, Curve). Our method achieves the highest IoU in most classes—especially Background, Car, and Person—and significantly outperforms others in mean IoU, demonstrating strong generalization and semantic consistency across categories.
Furthermore, Fig. 13 shows semantic segmentation visualizations on complex scenes. U2Fusion and CDDFuse exhibit blurred boundaries and texture loss, struggling with multiple targets and fine details. MaeFuse and SAGE improve coherence but still missegment small objects and complex backgrounds. Our method achieves clearer boundaries and more complete contours, demonstrating superior structural fidelity and semantic consistency.
@@ 5 DISCUSSION
Despite the state-of-the-art performance achieved by SGDFuse, there remain avenues for further exploration that constitute our future research directions. First, since SAM’s pre-training data is primarily based on visible light images, its generalization capability to the infrared domain has room for improvement. In the future, the precision of the semantic guidance can be further enhanced through domain-adaptive fine-tuning or by incorporating more specialized infrared segmentation models. Second, the iterative nature of the diffusion model limits its application in real-time scenarios. Future work will focus on model lightweighting, for instance, through knowledge distillation or by designing efficient samplers, to develop a version suitable for real-time applications and thereby broaden its operational boundaries.
@@ Acknowledgments
This research was supported by the National Natural Science Foundation of China (62301105, 61772319, 62272281, 62002200, 62202268), Shandong Natural Science Foundation of China (ZR2020QF012 and ZR2021MF068), Yantai science and technology innovation development plan(2022JCYJ031).

### v4

@@ 4.7.1 Object Detection
Figure 17 shows that fused images significantly improve detection accuracy over single-modality inputs. Our method outperforms mainstream fusion approaches in both “Person” and “Car” categories, demonstrating superior object perception and generalization. As shown in Figure 18, our method outperforms others in accurately localizing multiple targets by effectively preserving infrared thermal information and integrating visible details, thereby enhancing detection accuracy, robustness, and adaptability in complex scenes.
@@ 4.7.2 Semantic Segmentation
Table 9 shows segmentation results for six categories (Car, Person, Bike, Color Cone, Background, Curve). Our method achieves the highest IoU in most classes—especially Background, Car, and Person—and significantly outperforms others in mean IoU, demonstrating strong generalization and semantic consistency across categories.
Furthermore, Figure 19 shows semantic segmentation visualizations on complex scenes. U2Fusion and CDDFuse exhibit blurred boundaries and texture loss, struggling with multiple targets and fine details. MaeFuse and SAGE improve coherence but still missegment small objects and complex backgrounds. Our method achieves clearer boundaries
and more complete contours, demonstrating superior structural fidelity and semantic consistency.
@@ 5 DISCUSSION
Despite SGDFuse achieving SOTA performance, there remain broad avenues for future exploration. Our current ”formally simple yet effective” implementation should be regarded as one effective path, or a powerful baseline, within the SGG paradigm. Therefore, future exploration can proceed along several main directions: First, one could explore more ’complex’ or ’sophisticated’ guidance mechanisms. For instance, by designing a lightweight ’semantic encoder’ to replace simple concatenation, thereby encoding semantic information into more refined features; or by injecting multi-scale semantic guidance at each timestep of the diffusion process. Second, regarding guidance quality, given that SAM’s pre-training data is primarily based on visible light images, its generalization capability to the infrared domain—though proven effective in our experiments—still has room for improvement. In the future, the precision of the semantic guidance could be further enhanced through domain-adaptive fine-tuning. Third, future work will focus on model lightweighting, for instance, through knowledge distillation or by designing efficient samplers, to develop a version suitable for real-time applications and thereby broaden its operational boundaries.
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China under grant Nos. U24A20219, 62272281, 62576193, 62202268, and 62301105; special funds of Taishan Scholars Project of Shandong Province under grant Nos. tsqn202306274 and tsqn202507240; Natural Science Foundation of Shandong Province under grant Nos. ZR2025QC712, ZR2025QC695 and ZR2025MS985; Yantai science and technology innovation development plan(2022JCYJ031).